(a) Find the polynomial degree 3 in term of three linear factors if P(1) = P(-2) = 0,



P(3) = 200 and P(-1) = 8.



(b) Find the partial fraction for



12x³ + 1/P(x)

1
Expert's answer
2021-11-09T16:29:06-0500

a)

Let "P_3(x)=a(x-x_1)(x-x_2)(x-x_3)."

Given "x_1=1, x_2=-2." Then


"P_3(x)=a(x-1)(x+2)(x-x_3)"

"P(3)=200"


"P_3(3)=a(3-1)(3+2)(3-x_3)=200"

"a(3-x_3)=20"

"P(-1)=8"


"P_3(-1)=a(-1-1)(-1+2)(-1-x_3)=8"

"a(-1-x_3)=-4"

"\dfrac{3-x_3}{-1-x_3}=\dfrac{20}{-4}"

"3-x_3=5+5x_3"

"x_3=-\dfrac{1}{3}"

"a(-1-(-\dfrac{1}{3}))=-4"

"a=6"

"P_3(x)=6(x-1)(x+2)(x+\dfrac{1}{3})"

"=(6x+2)(x^2+x-2)"

"=6x^3+6x^2-12x+2x^2+2x-4"

Answer:

"P(x)=6x^3+8x^2-10x-4"



b)


"\dfrac{12x^3+1}{P(x)}=\dfrac{12x^3+1}{6x^3+8x^2-10x-4}"

"=\dfrac{A}{6}+\dfrac{B}{x-1}+\dfrac{C}{x+2}+\dfrac{D}{x+\dfrac{1}{3}}"

"A(x-1)(x+2)(x+\dfrac{1}{3})+6B(x+2)(x+\dfrac{1}{3})"

"+6C(x-1)(x+\dfrac{1}{3})+6D(x-1)(x+2)"

"=12x^3+1"

"x=1:6B(3)(\dfrac{4}{3})=12+1=>B=\dfrac{13}{24}"

"x=-2: 6C(-3)(-\dfrac{5}{3})=-96+1=>C=-\dfrac{19}{6}"

"x=-\dfrac{1}{3}: 6D(-\dfrac{4}{3})(\dfrac{5}{3})=-\dfrac{4}{9}+1=>D=-\dfrac{1}{24}"

"x=0:-\dfrac{2}{3}A+\dfrac{2}{3}(\dfrac{13}{4})-\dfrac{1}{3}(-19)-2(-\dfrac{1}{4})=1"

"A=12"

Answer:

"\dfrac{12x^3+1}{6x^3+8x^2-10x-4}=2+\dfrac{13}{24(x-1)}"

"-\dfrac{19}{6(x+2)}-\dfrac{1}{24(x+\dfrac{1}{3})}"